Express each of the following as a product of prime factors only in exponential form:  729 x 64

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The factors of 729 = 3 × 3 × 3 × 3 × 3 × 3 = 3^6
The factors of 64 = 2 × 2 × 2 × 2 × 2 × 2 = 2^6
Then,
= (3^6× 2^6)
= 3^6× 2^6

Thanks